How do I find the solution to the following system of conic sections???
(x^2/4) + y^2 = 1
y = x + 1
I know that the first equation is an ellipse and the second is a line and there are as many as two possible solutions to the system.

Answer by scott8148(6628)

substituting ___ (x^2 / 4) + (x + 1)^2 = 1

x^2 + 4x^2 + 8x + 4 = 4

5x^2 + 8x = 0 ___ x(5x + 8) = 0 ___ x = 0 and x = -8/5

y = (0) + 1 = 1

y = (-8/5) + 1 = -3/5

the solution points are ___ (0,1) and (-8/5,-3/5)